OpenStudy (jordans_openstudy):

Write an equation of a line in point slope form that has a slope of -2 and passes through (5,-1)

OpenStudy (jordans_openstudy):

A) y + 1 = -2(x – 5) B) y -5 = -2(x – 1) C) y – 1 = -2(x – 5) D) y – 5 = -2(x + 1)

OpenStudy (luigi0210):

Point slope: \(\large y-y_1=m(x-x_1)\) \((x_1, y_1)\) m=slope

OpenStudy (theyankee):

Plug your equation into point slope form where m is the slope.

OpenStudy (jordans_openstudy):

im not really sure I'm so bad at this

OpenStudy (theyankee):

Well, look at it this way. Plug your Y coordinate in as Y1 to get Y-5=m(x-x1) Then plug in your slope. Y-5=-2(x-x1) Then plug in your x coordinate to get Y-5=-2(x-1)

OpenStudy (jordans_openstudy):

hmm i think I'm kinda getting it now